OpenStudy (anonymous):

A 1000-kg car traveling at 66 m/s hits a 2000-kg car at rest and the two travel together after the collision. Calculate the speed with which they move.

OpenStudy (anonymous):

initial momentum = final momentum \[m_1v_1+m_2v_2=v_f(m_1+m_2)\]

OpenStudy (anonymous):

so 1000x2000 +2000x v2 =vf(3000) ???

OpenStudy (anonymous):

No, pair the masses and velocities. So 1000 kg goes with 66 m/s and 2000 kg goes with 0 m/s (since the car is at rest) \[(1000~kg)(66~m/s)+(2000~kg)(0~m/s)=v_f(2000~kg+1000~kg)\]

OpenStudy (anonymous):

so the only variable left if \(v_f\)

OpenStudy (anonymous):

so its 68000/3000 which equals 22.66 right ?

OpenStudy (anonymous):

2000*0=0, so 66,000/3,000 = 22 m/s even

OpenStudy (anonymous):

Thank you very much !!

OpenStudy (anonymous):

you're welcome